OpenStudy (anonymous):

The only force acting on a 2.3 kg body as it moves along the positive x axis has an x component Fx = - 6x N, where x is in meters. The velocity of the body at x = 3.0 m is 8.0 m/s. At what positive value of x will the body have a velocity of 5.0 m/s? I'm not sure how to approach this. So far I have written down KE=(1/2)mv^2, but I haven't the slightest idea where to go from here...

OpenStudy (anonymous):

\[F = ma\] -6x = (2.3)a -60x/23 = a \[Vf^2 = Vo^2 + 2ad\] 5^2 = 8^2 + 2(3)a 25 = 64 + 6a -39 = 6a -13/2 = a -60x/23 = -13/2 -60x = -299/2 -120x = -299 x = 2.49 (I think. I just kind of made this up and have never dealt with a problem like this before, but I think this might be the answer, and I'd be pleasantly surprised to find out so.)